# PO-Flow: Flow-based Generative Models for Sampling Potential Outcomes and Counterfactuals

### v2

@@ 5.3.1 Learning potential outcome densities
Table 3: KL divergence of learned potential outcome distributions across datasets. Results are averaged over 10-fold cross-validation.
ACIC 2016 ACIC 2018 IHDP IBM
PO-Flow $\mathbf{0.14_{\pm.04}}$ $\mathbf{0.08_{\pm.02}}$ $\mathbf{0.09_{\pm.03}}$ $\mathbf{0.08_{\pm.03}}$
GP-ITE $1.73_{\pm.34}$ $0.73_{\pm.19}$ $1.42_{\pm.30}$ $0.68_{\pm.17}$
Baselines: A natural baseline with density estimation capability is Gaussian Processes for interventional treatment effects, as used in [3, 26].
Results: PO-Flow achieves much lower KL divergence than GP baselines on the ACIC 2018, IHDP, and IBM datasets, indicating more accurate density estimation of potential outcomes.
This property enables PO-Flow to identify most likely potential outcomes based on their log-densities. Table 4 reports the RMSE of POs selected by PO-Flow with the highest log-density, illustrating one application of this property.
Table 4: Root mean squared error (RMSE) of estimated potential outcomes (POs) across different methods on the ACIC 2018, IHDP, and IBM datasets. For PO-Flow, results are based on the most likely PO selected from 100 simulated outcomes using log-density.
ACIC 2018 IHDP IBM
Compared with the averaged RMSE over all POs generated by PO-Flow in Table 1, selecting the most likely POs based on log-density yields smaller RMSEs, providing more confident and closer estimates.
5.3.2 Empirical distribution estimations
Evaluation Metrics: To enable comparison of estimated potential outcome distributions across methods, we include empirical distribution estimates for all baselines. These are evaluated using the Wasserstein-1 distance, defined as: $W_{1}(p_{\theta,Y|X,A}(y|x,a),p_{Y|X,A}(y|x,a))=\inf_{\gamma\in\Gamma(p_{\theta},p)}\mathbb{E}_{(\hat{y},y\sim\gamma)}[\|\hat{y}-y\|_{1}]$. The results are shown in Table 5.
Table 5: In-sample and out-of-sample Wasserstein distances across different methods on the ACIC 2016, ACIC 2018, IHDP, and IBM datasets. Results are averaged over 10-fold cross-validation.
ACIC 2016 ACIC 2018 IHDP IBM
$\hat{W}_{1}^{\text{in}}$ $\hat{W}_{1}^{\text{out}}$ $\hat{W}_{1}^{\text{in}}$ $\hat{W}_{1}^{\text{out}}$ $\hat{W}_{1}^{\text{in}}$ $\hat{W}_{1}^{\text{out}}$ $\hat{W}_{1}^{\text{in}}$ $\hat{W}_{1}^{\text{out}}$
PO-Flow $\mathbf{0.36_{\pm.11}}$ $\mathbf{0.42_{\pm.12}}$ $\mathbf{0.05_{\pm.03}}$ $\mathbf{0.12_{\pm.04}}$ $\mathbf{0.30_{\pm.05}}$ $\mathbf{0.41_{\pm.06}}$ $\mathbf{0.33_{\pm.07}}$ $\mathbf{0.44_{\pm.08}}$
[…]
INFs $0.59_{\pm.13}$ $67_{\pm.14} 0.0$$08_{\pm.03}$ $0.16_{\pm.05}$ $0.32_{\pm.08}$ $\mathbf{0.40_{\pm.09}}$ $0.36_{\pm.08}$ $\mathbf{0.45_{\pm.09}}$
S-learner $1.01_{\pm.13}$ $1.02_{\pm.20}$ $0.35_{\pm.13}$ $0.49_{\pm.20}$ $0.71_{\pm.14}$ $0.85_{\pm.19}$ $0.46_{\pm.11}$ $0.55_{\pm.14}$
T-learner $1.22_{\pm.15}$ $1.48_{\pm.18}$ $1.04_{\pm.13}$ $1.00_{\pm.30}$ $0.83_{\pm.20}$ $0.92_{\pm.28}$ $1.21_{\pm.23}$ $1.67_{\pm.32}$
CEVAE $0.62_{\pm.17}$ $0.70_{\pm.21}$ $0.88_{\pm.18}$ $1.02_{\pm.23}$ $0.65_{\pm.14}$ $0.76_{\pm.20}$ $0.38_{\pm.07}$ $\mathbf{0.45_{\pm.07}}$
CFR $0.68_{\pm.13}$ $0.73_{\pm.15}$ $0.88_{\pm.12}$ $0.90_{\pm.15}$ $0.50_{\pm.17}$ $0.56_{\pm.19}$ $0.92_{\pm.20}$ $0.93_{\pm.21}$
GANITE $0.74_{\pm.20}$ $0.86_{\pm.23}$ $0.89_{\pm.24}$ $1.18_{\pm.25}$ $0.71_{\pm.23}$ $0.82_{\pm.29}$ $1.09_{\pm.27}$ $1.34_{\pm.30}$
Understanding the distribution of potential outcomes (POs) is essential for uncertainty-aware treatment decisions, as outcomes for the same individual may vary substantially across treatments. PO-Flow provides not only point estimates but also full distributional insight, enabling practitioners to assess whether outcomes fall within desired ranges and to evaluate their likelihood (Proposition 4.7).
5.4 Counterfactual Outcome Estimation
Evaluation Metrics: We assess counterfactual predictions using the RMSE between model-estimated $\hat{y}_{\text{cf}}^{(1-a)}$ and ground-truth $y_{\text{cf}}^{(1-a)}$. Results are in Table 6.
Baselines: To our knowledge, very few existing methods are specifically designed for counterfactual outcome prediction conditioned on factuals in general settings—without explicitly modeling structural causal models or relying on temporal information. While methods like CFR include “counterfactual” in name, they target unobserved potential outcomes rather than factual-conditioned ones, though we still include them as valuable baselines. We also adapted DiffPO and GANITE for this task, as detailed in Appendix B.5.
Table 6: Root mean squared error (RMSE) of estimated counterfactual outcomes across different methods on the ACIC 2016, ACIC 2018, IHDP, and IBM datasets. Results are averaged over 10-fold cross-validation.
ACIC 2016 ACIC 2018 IHDP IBM
$\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$
PO-Flow $\mathbf{1.50_{\pm.15}}$ $\mathbf{1.53_{\pm.17}}$ $\mathbf{0.04_{\pm.01}}$ $\mathbf{0.05_{\pm.02}}$ $\mathbf{1.63_{\pm.09}}$ $1.69_{\pm.10}$ $\mathbf{0.04_{\pm.01}}$ $\mathbf{0.04_{\pm.02}}$
DiffPO $1.77_{\pm.20}$ $1.78_{\pm.20}$ $0.62_{\pm.14}$ $0.65_{\pm.15}$ $1.81_{\pm.22}$ $1.85_{\pm.23}$ $0.62_{\pm.13}$ $0.64_{\pm.16}$
CFR $15_{\pm.28} 3.3$$38_{\pm.30}$ $0.99_{\pm.05}$ $1.01_{\pm.07}$ $\mathbf{1.63_{\pm.21}}$ $\mathbf{1.65_{\pm.25}}$ $0.86_{\pm.18}$ $0.87_{\pm.16}$
GANITE $3.41_{\pm.52}$ $3.67_{\pm.55}$ $0.68_{\pm.13}$ $0.70_{\pm.14}$ $2.48_{\pm.53}$ $2.72_{\pm.59}$ $1.84_{\pm.30}$ $2.02_{\pm.37}$
5.5 Discussions and Comparisons
Training efficiency: PO-Flow exhibits high training efficiency, converging in approximately 300 gradient steps across multiple causal datasets (Appendix D.3), and demonstrates notable advantages in both training and sampling time compared to other baselines (Table 9).
Density estimation for principled uncertainty quantification: Our continuous normalizing flow (CNF) backbone enables density learning of potential outcomes, as detailed in Section 5.3.1. This property allows us to generate multiple potential outcomes for each individual and to identify the most likely ones based on their log-densities. The empirical evaluation of this procedure is presented in Table 4.
Counterfactual: We take a step further by exploring counterfactual outcomes, enabled by the smooth PO-Flow mapping that preserves factual information during encoding. Under (A1)–(A3), we include a supporting recovery result for this architecture (Proposition 12; Corollary 4.5). It is among the first to explore counterfactual outcomes in the potential outcomes framework without relying on temporal data or explicitly modeling structural causal models.
6 Conclusions
We introduced PO-Flow, a continuous normalizing flow framework for modeling potential and counterfactual outcomes in causal inference. PO-Flow provides a unified generative framework for potential outcome prediction, CATE estimation, and general-purpose counterfactual prediction in potential outcomes—without explicitly modeling structural causal models (SCMs) or relying on auxiliary information. To our knowledge, this constitutes one of the earliest attempts to address counterfactual outcome prediction in the potential outcomes literature, a crucial yet underexplored direction.
Theoretically, we include a supporting recovery result for the proposed architecture (Corollary 4.5). Empirically, PO-Flow attains state-of-the-art performance on benchmark datasets across standard causal tasks, while additionally supporting counterfactual outcome prediction and log-density estimation for principled uncertainty quantification. In addition, PO-Flow is both computationally efficient and robust compared to modern baselines. These results highlight PO-Flow as a unified and scalable solution for potential-outcomes–based causal tasks, supporting reliable decision-making in domains such as healthcare.
@@ Acknowledgment
D.I. acknowledges support from NSF (IIS-2212097), ARL (W911NF-2020-221), and ONR (N00014-23-C-1016). Y.X. acknowledges partial support from NSF DMS-2134037, CMMI-2112533, and the Coca-Cola Foundation. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsor(s).
@@ Appendix A Proofs
Proposition 12. (Encoded $Z$ as a function of the exogenous noise $U$) Let Assumption 4.1 hold. Assume that the exogenous noise $U\sim\mathrm{Unif}[0,1]$. Then there exists a continuously differentiable bijection function $\psi_{X}:\;\mathcal{U}\;\longrightarrow\;\mathcal{Z}$ that is independent of the treatment assignment $A$, i.e.,
@@
as a function that may depend on the action $A$. Therefore, our target becomes proving that $m_{X,A}(U)$ is functionally invariant to $A$.
By (A3) in Assumption 4.1, we have $Z\perp\!\!\!\perp A\ |\ X$. Therefore,
$p_{\theta,Z\mid X,A}(z)=p_{\theta,Z}(z).$ (23)
@@
Then based on (23), we have:
@@
Corollary 4.5 (Counterfactual recovery) Assume the conditions in Assumption 4.1 hold. Consider a factual sample $\bigl(Y^{(A)},X,A\bigr)$ given by the structural equation $Y^{(A)}=f_{X}\bigl(A,U\bigr)$, and let its encoded representation be $Z=\Phi_{\theta}(Y^{(A)};X,A)$. Apply the intervention $\mathrm{do}\bigl(A=\gamma\bigr)$. Assume that the true counterfactual is given by $Y_{\mathrm{CF}}^{(\gamma)}=f_{X}(\gamma,U)$.
@@
From Proposition 4.3, we have $Z=\;\Phi_{\theta}\bigl(Y^{(A)};\,X,A\bigr)\;=\Phi_{\theta}(f_{X}(A,U);\ X,A\bigr)\;=\;\psi_{X}\bigl(U\bigr)$ does not depend on $A$. Therefore, fixing $U$ as the same and plugging in $\mathrm{do}\bigl(A=\gamma\bigr)$, we have:
@@
Noting that the L.H.S. of (36) is exactly the counterfactual procedures (encoding-decoding) of PO-Flow, and by (33), the R.H.S. of (36) equals to $Y_{\mathrm{CF}}^{\gamma}$. Therefore,
@@
Proposition 4.7. Given a noise sample $z\sim q(\cdot)$ at $t=1$, the log-density of the resulting potential outcome at $t=0$, obtained via PO-Flow, is given by:
$\log p_{\theta,Y\mid X,A}\bigl(y(0)\mid x,a,z\bigr)=\log q(z)\;+\;\int_{0}^{1}\nabla_{y}\!\cdot\,v_{\theta}\bigl(y(t),t;\,x,a\bigr)\,dt.$
@@ Appendix B Experimental Details
B.1 Synthetic Causal Datasets
@@
The ACIC 2016 dataset contains 4,802 samples with 82 covariates each ($d_{X}=82$), ACIC 2018 has 1,000 samples with 177 covariates, IHDP includes 747 samples with 25 covariates, and the IBM dataset consists of 1,000 samples with 177 covariates. Each dataset provides three synthetic values per individual: the factual outcome $y^{(a)}$, the assigned treatment $a$, and the mean potential outcomes under both treatments, $\mu_{0}$ (for $a=0$) and $\mu_{1}$ (for $a=1$).
@@ 2. Action: intervene by changing the treatment assignment to 1A1-A;
B.2 Neural Network Structure and Implementation Information
@@
All experiments are conducted on an A100 GPU with 16 GB of memory. The training time for causal datasets depends on the dataset and batch sizes. For example, training on the ACIC 2018 dataset with a batch size of 256 typically takes 2–4 seconds per epoch. Log-density evaluation, as discussed in Section 4.3, involves divergence integration and requires significantly more time than other test metrics.
B.3 Runge-Kutta numerical integration method
@@
B.4 Hutchinson trace estimator
@@
B.5 Adapted Implementation Details for DiffPO and GANITE Counterfactuals
@@ Appendix C Comparisons to BGM settings
Bijective Generation Mechanisms (BGM). The BGM framework [43] shows that if the true structural mechanism $f$ is in the BGM class (bijective/strictly monotone in the exogenous noise), then any learned mechanism $\hat{f}$ that (i) is also in the BGM class and (ii) matches the observed distribution—i.e., for the same parents value $X$ and action $A$,
@@
PO-Flow. Our findings are specific to a particular methodology and architecture, as this is a methodology paper where we analyze continuous normalizing flows (CNFs). With the additional assumption (A3) specific to PO-Flow, Proposition 4.3 shows that the PO-Flow encoded latent $Z$ is bijective in the exogenous noise $U$. Consequently, Corollary 4.5 proves that the encode–decode procedure recovers the true counterfactual.
Relationships. Because of Proposition 4.3 (enabled by the CNF), PO-Flow implements a bijective-in-noise mechanism and is thus comparable to BGM in the sense that $\hat{f}$ is bijective in $U$. However, BGM additionally requires observational distribution matching to obtain class-level identifiability, whereas the proofs of Proposition 4.3 and Corollary 4.5 do not assume such matching.
Alternative route. The counterfactual recovery result (Corollary 4.5) can alternatively be established by imposing an additional assumption:
@@
Under (A1)–(A3), Proposition 4.3 establishes a continuously differentiable bijection $\psi_{X}:U\to Z$ (independent of $A$), so the induced mechanism $\hat{f}_{X}$ implemented by PO-Flow is bijective in the exogenous noise, i.e., PO-Flow lies in the BGM class. With the additional assumption (A4), we can therefore directly invoke the counterfactual recovery result of the BGM framework [43].

### v3

@@ VI-C Task 3: Learning Potential Outcome Densities
TABLE III: KL divergence of learned potential outcome densities (10-fold cross-validation). Parentheses indicate standard deviation.
Method ACIC16 ACIC18 IHDP IBM
PO-Flow 0.14(0.04) 0.08(0.02) 0.09(0.03) 0.08(0.03)
INF 0.22(0.09) 0.15(0.06) 0.14(0.07) 0.10(0.04)
GP-ITE 1.73(0.34) 0.73(0.19) 1.42(0.30) 0.68(0.17)
Likelihood evaluation also enables selecting the most likely potential outcomes under the learned density. Table IV reports the RMSE of PO-Flow (MAP), where predictions are chosen by maximizing the log-density.
TABLE IV: RMSE for PO-Flow variants (10-fold cross-validation). I/O denote in-/out-of-sample.
ACIC18 IHDP IBM
Method I O I O I O
PO-Flow 0.53 0.60 0.98 1.19 0.90 0.97
PO-Flow (MAP) 0.42 0.54 0.96 1.05 0.84 0.90
VI-D Task 4: Distributional Fit of Potential Outcomes
To compare estimated potential outcome distributions across methods, we evaluate empirical distribution fit using the Wasserstein-1 distance, following the standard definition
$\begin{split}&W_{1}\!\left(p_{\theta,Y\mid X,A}(\cdot\mid x,a),\,p_{Y\mid X,A}(\cdot\mid x,a)\right)\\ =&\inf_{\gamma\in\Gamma(p_{\theta},p)}\mathbb{E}_{(\hat{y},y)\sim\gamma}\bigl[\|\hat{y}-y\|_{1}\bigr],\end{split}$
where $\Gamma$ denotes the “coupling”: the joint distribution between a pair of random variables, whose marginal distributions match $p_{\theta}$ and $p$ respectively. Results are reported in Table V. PO-Flow consistently achieves strong performance, indicating accurate recovery of the full potential outcome distributions.
TABLE V: Wasserstein distances (standard deviations omitted, and they are reasonably small). I/O denote in-/out-of-sample.
ACIC16 ACIC18 IHDP IBM
Method I O I O I O I O
PO-Flow 0.36 0.42 0.05 0.12 0.30 0.41 0.33 0.44
DiffPO 0.47 0.51 0.04 0.13 0.80 0.88 0.69 0.75
INFs 0.59 0.67 0.08 0.16 0.32 0.40 0.36 0.45
S-learner 1.01 1.02 0.35 0.49 0.71 0.85 0.46 0.55
T-learner 1.22 1.48 1.04 1.00 0.83 0.92 1.21 1.67
CEVAE 0.62 0.70 0.88 1.02 0.65 0.76 0.38 0.45
CFR 0.68 0.73 0.88 0.90 0.50 0.56 0.92 0.93
GANITE 0.74 0.86 0.89 1.18 0.71 0.82 1.09 1.34
VI-E Task 5: Factual-Conditioned Counterfactual Outcome Estimation
We evaluate factual-conditioned counterfactual prediction using RMSE between the predicted counterfactual $\hat{y}_{\mathrm{cf}}^{(1-a)}$ and the ground-truth $y_{\mathrm{cf}}^{(1-a)}$. Results are shown in Table VI, which shows the proposed method achieves the best performance. While DiffPO and GANITE do not natively support counterfactual queries, we adapt them for evaluation. Details are provided in Appendix B.5 of [35].
TABLE VI: RMSE of estimated counterfactual outcomes (10-fold cross-validation). DiffPO and GANITE do not natively support this query and are adapted by us for evaluation. I/O denote in-/out-of-sample.
ACIC16 ACIC18 IHDP IBM
Method I O I O I O I O
PO-Flow 1.50 1.53 0.04 0.05 1.63 1.69 0.04 0.04
DiffPO 1.77 1.78 0.62 0.65 1.81 1.85 0.62 0.64
GANITE 3.41 3.67 0.68 0.70 2.48 2.72 1.84 2.02
Compared to DiffPO and INFs: PO-Flow supports factual-conditioned counterfactual prediction beyond standard potential-outcome modeling. While DiffPO and INFs primarily model interventional outcome distributions, PO-Flow performs counterfactual inference via an encoding-decoding procedure enabled by a smooth and information-preserving path of continuous normalizing flows. Although both baselines can be adapted to this task, they consistently underperform PO-Flow, likely due to INFs’ restrictive discrete-flow parameterization and DiffPO’s forward noising, which limit preservation of factual information.
VII Conclusions
We proposed PO-Flow, a continuous normalizing flow framework for causal inference that models potential outcomes and factual-conditioned counterfactuals within the potential outcomes framework. By mapping an observed factual outcome to a shared latent representation and decoding it under alternative treatments, PO-Flow enables individual-level factual–counterfactual mapping without requiring a parametric structural causal model. This supports individualized “what-if” analysis with likelihood-based uncertainty assessment, and the accompanying theory and experiments show that PO-Flow offers a unified and scalable approach to individualized causal reasoning.
@@ Acknowledgment
The work of D.W. and Y.X. are supported by NSF CMMI-2112533, and the Coca-Cola Foundation.
@@ Appendix A Proofs
Proposition V.4. (Encoded $Z$ as a function of the exogenous noise $U$) Let Assumptions III.1 and V.1 hold. Assume that the exogenous noise $U\sim\mathrm{Unif}[0,1]$. Then there exists a continuously differentiable bijection function $\psi_{X}:\;\mathcal{U}\;\longrightarrow\;\mathcal{Z}$ that is independent of the treatment assignment $A$, i.e.,
@@
as a function that may depend on the treatment $A$. Therefore, our target becomes proving that $m_{X,A}(U)$ is functionally invariant to $A$.
By (A3) in Assumption V.1, we have $Z\perp\!\!\!\perp A\ |\ X$. Therefore,
$p_{\theta,Z\mid X,A}(z)=p_{\theta,Z|X}(z).$ (19)
@@
Next, by combining (19) and (20), we have:
@@
Corollary V.6 (Counterfactual recovery) Assume the conditions in Assumptions III.1 and V.1 hold. Consider a factual sample $\bigl(Y^{(A)},X,A\bigr)$ given by the structural equation $Y^{(A)}=f_{X}\bigl(A,U\bigr)$, and let its encoded representation be $Z=\Phi_{\theta}(Y^{(A)};X,A)$. Apply the intervention $\mathrm{do}\bigl(A=\gamma\bigr)$. Assume that the true counterfactual is given by $Y_{\mathrm{CF}}^{(\gamma)}=f_{X}(\gamma,U)$.
@@
From Proposition V.4, we have $Z=\;\Phi_{\theta}\bigl(Y^{(A)};\,X,A\bigr)\;=\Phi_{\theta}(f_{X}(A,U);\ X,A\bigr)\;=\;\psi_{X}\bigl(U\bigr)$ does not depend on $A$. Therefore, fixing $U$ as the same and plugging in $\mathrm{do}\bigl(A=\gamma\bigr)$, we have:
@@
Noting that the L.H.S. of (32) is exactly the counterfactual procedures (encoding-decoding) of PO-Flow, and by (29), the R.H.S. of (32) equals $Y_{\mathrm{CF}}^{\gamma}$. Therefore,
@@
Proposition V.7. Given a noise sample $z\sim q(\cdot)$ at $t=1$, the log-density of the resulting potential outcome at $t=0$, obtained via PO-Flow, is given by:
$\log p_{\theta,Y\mid X,A}\bigl(y\mid x,a,z\bigr)=\log q(z)\;+\;\int_{0}^{1}\nabla_{y}\!\cdot\,v_{\theta}\bigl(y(t),t;\,x,a\bigr)\,dt.$
@@ Appendix B Experimental Details
B-A Synthetic Causal Datasets
@@
The ACIC 2016 dataset contains 4,802 samples with 82 covariates each ($d_{X}=82$), ACIC 2018 has 1,000 samples with 177 covariates, The IHDP dataset includes 747 samples with 25 covariates, and the IBM dataset consists of 1,000 samples with 177 covariates. Each dataset provides three synthetic values per individual: the factual outcome $y^{(a)}$, the assigned treatment $a$, and the mean potential outcomes under both treatments, $\mu_{0}$ (for $a=0$) and $\mu_{1}$ (for $a=1$).
@@ 2. Action: intervene by changing the treatment assignment to 1A1-A;
B-B Neural Network Structure and Implementation Information
@@
All experiments are conducted on an A100 GPU with 16 GB of memory. The training time for causal datasets depends on the dataset and batch sizes. For example, training on the ACIC 2018 dataset with a batch size of 256 typically takes 2–4 seconds per epoch. Log-density evaluation involves divergence integration and requires significantly more time than other test metrics.
B-C Runge-Kutta numerical integration method
@@
B-D Hutchinson trace estimator
@@
B-E Adapted Implementation Details for DiffPO and GANITE Counterfactuals
@@ Appendix C Comparisons to BGM settings
Bijective Generation Mechanisms (BGM). The BGM framework [25] shows that if the true structural mechanism $f$ is in the BGM class (bijective/strictly monotone in the exogenous noise), then any learned mechanism $\hat{f}$ that (i) is also in the BGM class and (ii) matches the observed distribution—i.e., for the same parents value $X$ and treatment $A$,
@@
PO-Flow. Our findings are specific to a particular methodology and architecture, as this is a methodology paper where we analyze continuous normalizing flows (CNFs). With the additional assumption (A3) specific to PO-Flow, Proposition V.4 shows that the PO-Flow encoded latent $Z$ is bijective in the exogenous noise $U$. Consequently, Corollary V.6 proves that the encode–decode procedure recovers the true counterfactual.
Relationships. Because of Proposition V.4 (enabled by the CNF), PO-Flow implements a bijective-in-noise mechanism and is thus comparable to BGM in the sense that $\hat{f}$ is bijective in $U$. However, BGM additionally requires observational distribution matching to obtain class-level identifiability, whereas the proofs of Proposition V.4 and Corollary V.6 do not assume such matching.
Alternative route. The counterfactual recovery result (Corollary V.6) can alternatively be established by imposing an additional assumption:
@@
Under (A1)–(A3), Proposition V.4 establishes a continuously differentiable bijection $\psi_{X}:U\to Z$ (independent of $A$), so the induced mechanism $\hat{f}_{X}$ implemented by PO-Flow is bijective in the exogenous noise, i.e., PO-Flow lies in the BGM class. With the additional assumption (A4), we can therefore directly invoke the counterfactual recovery result of the BGM framework [25].